OpenStudy (anonymous):

Mike wants to make meatloaf. His recipe uses a total of 5 pounds of meat. Q: If he uses a 3 to 1 ratio of beef to pork will he use? Enter you answer as a mixed number in the simplest terms. PLEASE HELP!

OpenStudy (anonymous):

Is the question the amounts of beef and pork meat?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Let a be beef and b pork. Thus: a + b = 5 a = 3b Solve the simultaneous equations.

OpenStudy (anonymous):

5/4 , 15/4

OpenStudy (anonymous):

i got O o

OpenStudy (anonymous):

Zuige, how did you get 3b? It doesn't make sense to me.

OpenStudy (anonymous):

ratio 3:1 3/1 = a/b a = 3b

OpenStudy (anonymous):

Ishaan94, you said you got 0? Correct?

OpenStudy (anonymous):

The answers are: Beef = 3 3/4 pounds (mixed) Pork = 1 1/4 pounds (mixed)

OpenStudy (anonymous):

Oh..okay. I understand it. Thanks, Zuige!